# Cross-composition Feature Disentanglement for Compositional Zero-shot Learning

### v1

@@ Introduction
Cross-composition Information in Adapters
@@
We further validate the effectiveness of introducing primitive-sharing compositions for each target composition by deleting the compositional graph in L-Adapters and/or auxiliary compositions in V-Adapters. Concretely, we merely keep the prompts’ token embeddings trainable in L-Adapters; and/or replace the two auxiliary images with the target image itself in V-Adapters, which thus turns into the self-attention on the input image, but still projects the primitive features into different subspaces with different attention networks. The results on MIT-States are shown in the last three lines of Table 2. The performance drop indicates the superiority of introducing primitive-sharing compositions in both text and image encoders. In particular, we find that V-Adapters without primitive-sharing compositions even perform worse than removing the whole V-Adapters, illustrating that these compositions play a considerable role in constraining the learning of primitive features in different subspaces. In contrast, L-Adapters without neighboring (primitive-sharing) compositions perform better than removing the whole L-Adapters. However, the performance gap is slight, which may be attributed to that we also set the embeddings of tokens in the prompts tunable, the cross-composition primitive features can be implicitly captured by optimizing the token embeddings with multiple samples.
Ablation Studies
We also ablate our model w.r.t the insertion location and depth of Adapters on MIT-States in closed world.
Insertion Location of Adapters
@@
There are two choices for inserting adapters into a transformer block, i.e., inside or outside. Therefore, we experiment with four configurations where every L-Adapter and V-Adapter are added inside or outside a transformer block, as shown in Table 3. Notably, the inside insertion includes adding after the self-attention and feed-forward layers in a block, i.e., the positions ③ and ② in Figure 2(c). From Table 3, it can be seen that adding V-Adapters outside transformer layers often achieves better performance no matter where L-Adapters are located, while there is no significant difference when shifting L-Adapters at different positions. This may be because when adding a V-Adapter inside the transform layer, there is no nonlinear transformation like Feed-Forward layer between two attention operations for extracting more informative features. Regarding the better performance with inside L-Adapters and outside V-Adapters, we finally apply this configuration to three benchmarks.
@@ I O I O S U H AUC
Table 3: Ablation Study on Adapters’ Insertion Locations, i.e., inside (I) or outside (O) one transformer block.
Insertion Depth of Adapters
@@
We further show the performance change when we increase the number of transformer blocks with trainable adapters in Figure 3. As mentioned earlier, we start from the top transformer layers. It is clear that the best performance is achieved with the last 3 layers, while the last 6 layers may overfit the training data, with less generalization knowledge from our adapters and CLIP.
Conclusions and Outlook
In this study, we proposed DCDA, a novel CLIP-based CZSL method, which includes dual trainable adapters (i.e., L-Adapters and V-Adapters) inserted into the frozen text and image encoders of CLIP for learning textual and visual cross-composition disentangled representations for attribute and object primitives in CZSL. Specifically, L-Adapters are built upon a compositional graph with features of different composition neighbors aggregated to their centric attributes or objects, while V-Adapters are built upon the cross-attention over multiple primitive-sharing compositions with a sampling strategy guided by two primitive-relevance graphs. Unlike the disentangled primitive features learned previously, we derive more discriminative and more general primitive features, with promising results achieved on three popular benchmarks. Moreover, in our experiments, we find that the performance gain is mostly from V-Adapters, illustrating the serious entanglement of primitive features in the image data. While the entanglement of textual primitive features is also noticeable as the incorporation of L-Adapters leads to better performance. In the future, it is expected to apply our dual adapters in other vision-language tasks such as visual question answering (Chen et al. 2021) and image captioning (Vinyals et al. 2015).
Appendix A Appendix
A. Supplementary Methodology Details
Computing the Sampling Probability
To sample the top-$n$ most and least relevant compositions according to the top-$n$ maximum and minimum primitive relevance scores, we take two different ways to compute the sampling probabilities. More specifically, the probability for selecting the top-$n$ most relevant compositions, e.g., $(a,o^{\prime}_{i})$, is computed as:
$P(o^{\prime}=o^{\prime}_{i})=\frac{A^{\text{obj}}_{o,o^{\prime}_{i}}}{\sum_{o^{\prime}_{k}\in\{o^{\prime}\}}A^{\text{obj}}_{o,o^{\prime}_{k}}},$
where $A^{\text{obj}}_{o,o^{\prime}_{i}}$ means the relevance score between objects $o^{\prime}_{i}$ and $o$, while the probability for selecting the top-$n$ least relevant compositions, e.g., $(a,o^{\prime}_{j})$, is computed as:
$P(o^{\prime}=o^{\prime}_{j})=-\frac{A^{\text{obj}}_{o,o^{\prime}_{j}}}{\sum_{o^{\prime}_{k}\in\{o^{\prime}\}}A^{\text{obj}}_{o,o^{\prime}_{k}}}.$
The difference is that the $n$ normalized minimum scores are multiplied by $-1$ to ensure that the top-1 least relevant composition has the highest probability of being sampled among the $n$ least relevant compositions. Consider the example in Figure 2(b), red wine, red rose and red leaf are top-3 least relevant compositions for red tomato, with top-3 normalized minimum object-relevance scores: $0.1,0.2$ and $0.2$, respectively, while red wine is the top-1 least relevant composition with the highest sampling probability among these three compositions.
B. Supplementary Experiment Details
Dataset
MIT-States contains 53,753 real-world images, annotated by a variety of classes with 245 objects and their 115 attributes in the general domain. In closed world, it provides 1,962 compositions in total, 1,262 of which are seen used for training, and 700 are unseen with 300 for validation and 400 for testing. UT-Zappos is a more domain-specific dataset, containing 50,025 images of shoes paired with their material attributes. In total, it has 16 attributes and 12 objects, yielding 83 seen and 33 unseen compositions under the closed world setting. C-GQA, derived from of Stanford GQA dataset (Hudson and Manning 2019), is the most extensive dataset for CZSL, containing 7,767 compositions (5,592/2,175 as seen/unseen), 413 attribute classes, 674 object classes, and 39,298 images in total. Table 4 summarizes the statistics of these three datasets in the closed world setting, the open world has the same set of testing images, i.e., $\mathcal{X}_{te}$, but larger candidate label set, i.e., all possible compositions $\mathcal{C}$ obtained by the Cartesian product of $\mathcal{A}$ and $\mathcal{O}$.
@@ B. Supplementary Experiment Details
Evaluation Metrics
We compute the prediction accuracy for recognizing seen and unseen compositions, i.e., the generalized CZSL, in both closed world and open world scenarios. Specifically, due to the inherent bias towards seen classes, we follow the current standard (Purushwalkam et al. 2019; Nayak, Yu, and Bach 2023) to add a scalar bias to the prediction scores of unseen classes and vary the bias from $\infty$- ∞ to +$\infty$+ ∞ to get a seen-unseen accuracy curve, which indicates the seen accuracy on the x-axis and unseen accuracy on the y-axis. Correspondingly, we can report the best seen accuracy (S), where the bias is set to $\infty$- ∞ and the models only predict on the seen labels, and report the best unseen accuracy (U), where the bias is set to +$\infty$+ ∞ and the models only predict on the unseen labels. Also, we calculate the best harmonic mean (H), where a harmonic mean value is first computed for each point on the curve to balance the seen accuracy ($acc_{s}$) and unseen accuracy ($acc_{u}$) as $(2\times acc_{s}\times acc_{u})/(acc_{s}+acc_{u})$, and then the highest value across all the selected points is reported. Finally, we compute the Area Under the accuracy Curve (AUC) as a comprehensive metric.
Experiments on C-GQA
To deal with the low-resolution and small-size images in C-GQA, we tried to add more trainable adapters in CLIP’s image encoder or fully fine-tune the whole image encoder, to adapt CLIP to this kind of image. More specifically, we add a downsample-upsample style adapter, which is similar to the adapter used in CAILA and Troika, after the self-attention layer and feed-forward layer in each vision transformer block (i.e., the positions ③ and ② in Figure 2(c)), except for the last three transformer blocks where our V-Adapters have already been there. In addition, we also try to freeze the whole image encoder to fully update its parameters, where our V-Adapters are added in the last three layers for feature disentanglement. The results are presented in the second and third lines of Table 5, respectively. Moreover, since a single A100 GPU with 40GB memory can not afford these extra adapters or fully fine-tuned parameters with CLIP-large that contains 24 vision transformer (ViT) layers, we instead use CLIP-base with ViT-B/32 as its image encoder, and re-run baselines for a fair comparison. Regarding that some baselines with CLIP-large still perform worse than our methods with CLIP-base, we omit re-implementing them to save computation costs.
@@
Table 5: Ablation on adding more trainable parameters in the vision encoder on C-GQA in closed world. All the results are tested with the “[PRG+N]” sampling method.
C. Supplementary Case Studies
@@
Figure 4: $t$-SNE visualizations of disentangled object representations of images in the test set of MIT-States, learned by CAILA and our DCDA. Solid and hollow circles represent images of seen and unseen compositions, respectively. Best viewed in color.

### v2

@@ 4.2 Main Results
4.3 Effectiveness of Adapters
The whole Adapters. We evaluate the contribution of L-Adapter and V-Adapter by analyzing the performance drop when one of them is removed. Notably, when all L-Adapters (reps. V-Adapters) are removed, the text (resp. image) encoder will act like CLIP’s default frozen encoders to output an entangled representation for each input prompt (resp. image). We conduct experiments on MIT-States dataset under the closed world setting, the results are shown in the second and third lines of Table 2. We can see that the performance both declines when L-Adapters or V-Adapters are removed, indicating that they two both have a positive contribution to the DCDA model and are complementary to each other. We also observe that the performance decrease of removing V-Adapters is greater than that of removing L-Adapters, which is consistent with our statement: textual primitive features are less entangled than visual ones, and the independent textual primitive features can still be captured by setting individual primitive prompts.
@@ 4.3 Effectiveness of Adapters
Cross-composition Information in Adapters. We further validate the effectiveness of introducing primitive-sharing compositions for each target composition by deleting the compositional graph in L-Adapters and/or auxiliary compositions in V-Adapters. Concretely, we merely keep the prompts’ token embeddings trainable in L-Adapters; and/or replace the two auxiliary images with the target image itself in V-Adapters, which thus turns into the self-attention on the input image, but still projects the primitive features into different subspaces with different attention networks. The results on MIT-States are shown in the last three lines of Table 2. The performance drop indicates the superiority of introducing primitive-sharing compositions in both text and image encoders. In particular, we find that V-Adapters without primitive-sharing compositions even perform worse than removing the whole V-Adapters, illustrating that these compositions play a considerable role in constraining the learning of primitive features in different subspaces. In contrast, L-Adapters without neighboring (primitive-sharing) compositions perform better than removing the whole L-Adapters. However, the performance gap is slight, which may be attributed to that we also set the embeddings of tokens in the prompts tunable, the cross-composition primitive features can be implicitly captured by optimizing the token embeddings with multiple samples.
4.4 Ablation Studies
@@
Insertion Location of Adapters. There are two choices for inserting adapters into a transformer block, i.e., inside or outside. Therefore, we experiment with four configurations where every L-Adapter and V-Adapter are added inside or outside a transformer block, as shown in Table 3. Notably, the inside insertion includes adding after the self-attention and feed-forward layers in a block, i.e., the positions ③ and ② in Figure 3(c). From Table 3, it can be seen that adding V-Adapters outside transformer layers often achieves better performance no matter where L-Adapters are located, while there is no significant difference when shifting L-Adapters at different positions. This may be because when adding a V-Adapter inside the transform layer, there is no nonlinear transformation like Feed-Forward layer between two attention operations for extracting more informative features. Regarding the better performance with inside L-Adapters and outside V-Adapters, we finally apply this configuration to three benchmarks.
@@ I O I O S U H AUC
Table 3: Ablation Study on Adapters’ Insertion Locations – inside (I) or outside (O) one transformer block.
@@
Insertion Depth of Adapters. We further show the performance change when we increase the number of transformer blocks with trainable adapters in Figure 4. As mentioned earlier, we start from the top transformer layers. It is clear that the best performance is achieved with the last 3 layers, while the last 6 layers may overfit the training data, with less generalization knowledge from our adapters.
5 Conclusions and Outlook
We present DCDA, a graph-guided framework that enhances CLIP for compositional zero-shot learning through dual adapters. Our key innovations include: 1) L-Adapters that aggregate textual primitives via compositional graph propagation, addressing label-side entanglement; 2) V-Adapters that extract invariant visual patterns through cross-attention and primitive-relevance guided sampling, effectively mitigating image feature entanglement. Experiments across three benchmarks validate that DCDA learns more discriminative and generalizable primitive representations. This underscores the critical role of visual disentanglement in CZSL, as visual primitives exhibit stronger composition-dependent entanglement than textual counterparts. In the future, it is expected to apply our dual adapters in other vision-language tasks and develop dynamic graph construction mechanisms to handle open-vocabulary primitive discovery.
Limitations
While our proposed DCDA demonstrates strong performance in compositional zero-shot learning (CZSL), several limitations warrant discussion as follows: (1) the computational overhead of constructing and updating the compositional graph grows with the scale of attributes and objects, which may pose challenges for applications requiring real-time inference on large combinatorial spaces. (2) while our primitive-relevance guided sampling mitigates data imbalance, extreme long-tailed distributions of attributes or objects (e.g., rare primitives with few compositions) may still lead to suboptimal disentanglement due to insufficient cross-composition supervision. Addressing these limitations could further enhance the robustness and scalability of our framework.
Acknowledgments
This work is partly funded by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ24F020034, and the Primary R&D Plan of Zhejiang under Grant No. 2023C03198. Jiaoyan Chen is mainly supported by the EPSRC project OntoEm (EP/Y017706/1). Xiang Chen is mainly supported by the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.1015-YAH24096), and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics.
Appendix A Supplementary Methodology Details
@@
Table 5: Ablation on adding more trainable parameters in the vision encoder on C-GQA in closed world. All results are tested with the “[PRG+N]” sampling method.
A.1 Computing the Sampling Probability
To sample the top-$n$ most and least relevant compositions according to the top-$n$ maximum and minimum primitive relevance scores, we take two ways to compute the sampling probabilities. More specifically, the probability for selecting the top-$n$ most relevant compositions, e.g., $(a,o^{\prime}_{i})$, and the probability for selecting the top-$n$ least relevant compositions, e.g., $(a,o^{\prime}_{j})$, are computed as, respectively:
$P(o^{\prime}=o^{\prime}_{i})=\frac{A^{\text{obj}}_{o,o^{\prime}_{i}}}{\sum_{o^{\prime}_{k}\in\{o^{\prime}\}}A^{\text{obj}}_{o,o^{\prime}_{k}}},P(o^{\prime}=o^{\prime}_{j})=-\frac{A^{\text{obj}}_{o,o^{\prime}_{j}}}{\sum_{o^{\prime}_{k}\in\{o^{\prime}\}}A^{\text{obj}}_{o,o^{\prime}_{k}}},$
where $A^{\text{obj}}_{o,o^{\prime}_{i}}$ means the relevance score between objects $o^{\prime}_{i}$ and $o$, the same applies to the others. The difference is that the normalized minimum scores are multiplied by $-1$ to ensure that the top-1 least relevant composition has the highest probability of being sampled among the $n$ least relevant compositions. Consider the example in Figure 3(b), red wine, red rose and red leaf are top-3 least relevant compositions for red tomato, with top-3 normalized minimum object-relevance scores: $0.1,0.2$ and $0.2$, respectively, while red wine is the top-1 least relevant composition with the highest sampling probability among these three compositions.
Appendix B Supplementary Experiment Details
B.1 Dataset
MIT-States contains 53,753 real-world images, annotated by a variety of classes with 245 objects and their 115 attributes in the general domain. In closed world, it provides 1,962 compositions in total, 1,262 of which are seen used for training, and 700 are unseen with 300 for validation and 400 for testing. UT-Zappos is a more domain-specific dataset, containing 50,025 images of shoes paired with their material attributes. In total, it has 16 attributes and 12 objects, yielding 83 seen and 33 unseen compositions under the closed world setting. C-GQA, derived from of Stanford GQA dataset Hudson and Manning (2019), is the most extensive dataset for CZSL, containing 7,767 compositions (5,592/2,175 as seen/unseen), 413 attribute classes, 674 object classes, and 39,298 images in total. Table 4 summarizes the dataset statistics in the closed world setting, the open world has the same set of testing images, i.e., $\mathcal{X}_{te}$, but larger candidate label set, i.e., all possible compositions $\mathcal{C}$ obtained by the Cartesian product of $\mathcal{A}$ and $\mathcal{O}$.
B.2 Evaluation Metrics
We compute the prediction accuracy for recognizing seen and unseen compositions, i.e., the generalized CZSL, in both closed world and open world scenarios. Specifically, due to the inherent bias towards seen classes, we follow the current standard Purushwalkam et al. (2019); Nayak et al. (2023) to add a scalar bias to the prediction scores of unseen classes and vary the bias from $\infty$- ∞ to +$\infty$+ ∞ to get a seen-unseen accuracy curve, which indicates the seen accuracy on the x-axis and unseen accuracy on the y-axis. As a result, we report the best seen accuracy (S), where the bias is set to $\infty$- ∞ and the models only predict on the seen labels, and report the best unseen accuracy (U), where the bias is set to +$\infty$+ ∞ and the models only predict on the unseen labels. We also calculate the best harmonic mean (H), where a harmonic mean value is first computed for each point on the curve to balance the seen accuracy ($acc_{s}$) and unseen accuracy ($acc_{u}$) as $(2\times acc_{s}\times acc_{u})/(acc_{s}+acc_{u})$, and then the highest value across all the selected points is reported. Finally, we compute the Area Under the accuracy Curve (AUC) as a comprehensive metric.
B.3 Implementation Details
We implement our models with PyTorch and use Adam as the optimizer, with the learning rate set to 5e-5, 5e-5, 1e-5, and the batch size set to 32, 32, 16 for MIT-States, UT-Zappos, C-GQA, respectively, the weight decay is set to 5e-5 for all datasets. The GNN module is implemented as GCN with $K=2$. $n$ is set to 5 for selecting the representative auxiliary compositions. The initialized values of $\alpha,\beta,\gamma$ are all set to 1, and then optimized together with other parameters. The optimal hyperparameter configurations are determined using AUC on the validation set. All the experiments are run on a single NVIDIA Tesla A100 GPU with 40GB memory.
B.4 Experiments on C-GQA
To deal with the low-resolution and small-size images in C-GQA, we tried to add more trainable parameters, such as inserting more adapters in CLIP’s image encoder or fully fine-tuning the whole image encoder, to adapt CLIP to this kind of image. More specifically, we add a downsample-upsample style adapter, which is similar to the adapter used in CAILA and Troika, after the self-attention layer and feed-forward layer in each vision transformer block (i.e., the positions ③ and ② in Figure 3(c)), except for the last three transformer blocks where our V-Adapters have already been there. In addition, we also try to unfreeze the whole image encoder to fully update its parameters, where our V-Adapters are added in the last three layers for feature disentanglement. The results are presented in the second and third lines of Table 5, respectively. Moreover, since a single A100 GPU with 40GB memory cannot afford these extra adapters or fully fine-tuned parameters with CLIP-large that contains 24 vision transformer (ViT) layers, we instead use CLIP-base with ViT-B/32 as its image encoder, and re-run baselines for a fair comparison. Regarding that some baselines with CLIP-large still perform worse than our methods with CLIP-base, we omit re-implementing them to save computation costs.
Appendix C Supplementary Case Studies
@@
Figure 5: $t$-SNE visualizations of disentangled object representations of images in the test set of MIT-States, learned by CAILA Zheng et al. (2024) and our DCDA. Solid and hollow circles represent images of seen and unseen compositions, respectively. Best viewed in color.